Chemists (and Astronomers!) like to express length
Answer #1

B)
r = 1*10^-10 m
q1=q2 = 1.6*10^-19 C
U = K*q1*q2/r
     = (9*10^9)*(1.6*10^-19)*(1.6*10^-19) / (1*10^-10)
     = 2.304*10^-18 J
    = (2.304*10^-18)/(1.6*10^-19)
    = 14.4 eV
Answer:
2.304*10^-18 J
14.4 eV

C)
rho = 3.5*10^-5 ohm-m
r = 1*10^-3 m
A = pi*r^2
    = pi*(1*10^-3)^2
    = 3.14*10^-6 m^2
R = 4000 ohm

R = rho*L/A
4000 = 3.5*10^-5 * L/ (3.14*10^-6)
L = 359 m
Answer: Length is 359 m

E = V/L
    = 5/359
     = 0.014 V/m
Answer: Electric field is 0.014 V/m

I am allowed to answer only 1 question at a time